Good day tutors of algebra.com
I dunno if I posted to the right category but here goes
Jack starts walking at 4kph . After a time , he hitches a ride with an old truck that travels 30kph . In 10 hours of walking and riding , he covers 157km . How long did he walk and how long did he ride?
I'm so confuse with this question because the distance was combined
thanks in advance

Answer by mananth(16949)

let him walk x km
so he trucked 157-x km
time walked = x/4
time trucked = (157-x)/30
Total time = 10 hours
x/4 + (157-x)/30 =10
LCD = 60
multiply the equation by LCD
15x+(157-x)*2 = 600
15x+314-2x=600
13x=600-314
13x=286
x=22 the distance walked
speed walked = 4 kph
time walked = 22/4 = 5.5 hours
m.ananth@hotmail.ca